OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12 A. y = -12 - 4/3 x B. y = -4 - 4/3 x C. y = -12 + 4x D. y = -4 + 4x

OpenStudy (anonymous):

4x + 3y = -12 => 3y = -12 - 4x => y = -12 -4/3 x Therefore, (A) y = -12 - 4/3 x is the correct option.

OpenStudy (anonymous):

4x + 3y = -12 \[ 4x \color{red}{-4x}+ 3y = -12 \color{red}{-4x} \\ \\ 3y=-12-4x \] Now divide both sides by 3

OpenStudy (anonymous):

ok. So y=-12-4/3x

OpenStudy (anonymous):

Thank you!